OpenStudy (anonymous):

PLEASE HELP AGAIN. U,U Quadrilateral ABCD is a kite. if AD = AB, Find BC. (the answers given) 7 17 25 10

OpenStudy (anonymous):



OpenStudy (anonymous):

i dont have the numbers to AD OR AB so heres where im stuck..

OpenStudy (mathstudent55):

What do you know about the lengths of the sides of a kite?

OpenStudy (anonymous):

just whats on the picture i drew. everything else isnt given.

OpenStudy (anonymous):

i dont know how im supposed to get either of those letters

OpenStudy (mathstudent55):

Ok, I understand that. I don't mean this kite. I mean what do you know about kites in general. What is the definition of a kite?

OpenStudy (anonymous):

a diformed square? lol. i havent had a kite in years..

OpenStudy (anonymous):

4 sides?

OpenStudy (anonymous):

in this diagram, if it is a kite, then, AB=AD and DC=BC. hence, 3x-4=2x+3=>x=7.

OpenStudy (anonymous):

ohh

OpenStudy (mathstudent55):

Ok. A kite does have 4 sides. It's a quadrilateral, but it's a special quadrilateral that has two sides congruent and the other two sides congruent. All 4 sides are not congruent. It has two pairs of congruent sides. 

OpenStudy (anonymous):

so its 7 right?

OpenStudy (anonymous):

both sides are congruent to each other. so there different equations?

OpenStudy (anonymous):

making it 3x-4=2x+3=

OpenStudy (anonymous):

my calculator says 7.

OpenStudy (mathstudent55):

Here you are told that AD = AB to let you know that pair of sides is congruent. Then you can conclude that CD and CB are also congruent.

OpenStudy (mathstudent55):

2x + 3 = 3x - 4 Subtract 3x from both sides: -x + 3 = -4 Subtract 3 from both sides: -x = -7 Divide both sides by -1 x = 7

OpenStudy (mathstudent55):

x = 7 is still not the final answer. BC = 3x - 4, so now you replace x with 7 to get BC